Determine your mass in kilograms. If you were to support that mass on one big toe, with an area of 2.000 cm2 (that is, 0.00020 m2), what would be the amount of pressure on that toe?
1
Expert's answer
2018-05-08T11:00:08-0400
Pressure is defined as force per unit area.
Pressure=Force/Area=F/A
For an object sitting on a surface, the force pressing on the surface is the weight of the object
F=mg
where g=9.8 m/s2.

Suppose your mass is
m=45 kg.
So, the pressure is
P=mg/A=((45)(9.8))/0.00020=2205000 Pa